OpenStudy (anonymous):

Multiply or divide as indicated. Simplify your answer, if possible. b^3 x b ^-3

OpenStudy (anonymous):

1

OpenStudy (anonymous):

b^3 x b ^-3=b^3 x b ^-3=b^3-3 =b^0=1

OpenStudy (anonymous):

just rem any number that has a power of zero =1

OpenStudy (sriram):

=b^3 / b^3 =1

OpenStudy (anonymous):

3^0=1,, 21^0=1 etc...